# Assessing the Uncertainty and Robustness of Object Detection Models for Detecting Stickers on Laptops

### v1

@@ Abstract
Refurbishing laptops extends their lives while contributing to reducing electronic waste, which promotes building a sustainable future. To this end, the Danish Technological Institute (DTI) focuses on the research and development of several applications, including laptop refurbishing. This has several steps, including cleaning, which involves identifying and removing stickers from laptop surfaces. DTI trained six sticker detection models (SDMs) based on open-source object detection models to identify such stickers precisely so these stickers can be removed automatically. However, given the diversity in types of stickers (e.g., shapes, colors, locations), identification of the stickers is highly uncertain, thereby requiring explicit quantification of uncertainty associated with the identified stickers. Such uncertainty quantification can help reduce risks in removing stickers, which, for example, could otherwise result in damaging laptop surfaces. For uncertainty quantification, we adopted the Monte Carlo Dropout method to evaluate the six SDMs from DTI using three datasets: the original image dataset from DTI and two datasets generated with vision language models, i.e., DALL-E-3 and Stable Diffusion-3. In addition, we presented novel robustness metrics concerning detection accuracy and uncertainty to assess the robustness of the SDMs based on adversarial datasets generated from the three datasets using a dense adversary method. Our evaluation results show that different SDMs perform differently regarding different metrics. Based on the results, we provide SDM selection guidelines and lessons learned from various perspectives.
Keywords Uncertainty Quantification  $\cdot$ Robustness Evaluation  $\cdot$ Object Detection  $\cdot$ Deep Neural Network  $\cdot$ Empirical Evaluation
@@ 1 Introduction
The European Union’s Circular Economy Action Plan (CEAP) highlights the need for sustainable operations to promote circular economy processes and encourage sustainable consumption [1]. One essential activity is refurbishing electronic devices (e.g., laptops) to extend their lives, reduce electronic waste, and provide affordable options for consumers. A critical and time-consuming step in refurbishment is removing stickers from the laptop by first identifying stickers and their locations on the laptops. Manual cleaning is time-consuming and faces challenges in finding enough workers, and current automation solutions are not built for this amount of variation, limiting sustainability and scalability. Thus, novel solutions to transition from manual labor to automated processes are paramount. Robotics offers a promising solution to simplify and scale up this process, increasing efficiency and reducing labor costs.
@@
The Danish Technological Institute (DTI) develops, applies, and transfers technology to industry and society. One leading area that DTI focuses on is laptop refurbishing automatically. Targeting the cleaning process in laptop refurbishment, DTI built several deep neural networks (DNN)–based sticker detection models (SDMs) for automatic sticker detection, which is the basis for successful automatic sticker removal. The SDMs are built on open-source object detection DNNs and trained using a sticker detection dataset specially designed by DTI. DNN has revolutionized various domains, such as computer vision [2] and autonomous systems [3], by providing powerful tools for predictive modeling and decision-making. However, due to inappropriate model architecture or insufficient training data, the design and training process may introduce uncertainties into the SDMs, making them vulnerable under certain conditions, such as adversarial attacks, noisy data, or unforeseen input distributions [4]. This vulnerability highlights the need to quantify uncertainty and evaluate the robustness of the SDMs, which is crucial for trustworthy sticker removal, as incorrect detection may damage the laptop surface or incomplete sticker removal.
Various approaches have been proposed to quantify DNN uncertainty. For example, Bayesian Neural Network (BNN) [5] provides a probabilistic framework for uncertainty quantification (UQ) through Bayesian inference, but it requires computationally complex posterior calculations. Monte Carlo Dropout (MC-Dropout) [6] is proposed as a Bayesian approximation based on dropout, offering a practical way to quantify prediction uncertainty by measuring the variability among multiple predictions. Regarding the robustness evaluation of DNNs, a typical technique is to evaluate the robustness under various adversarial attacks, referred to as adversarial robustness [7]. For instance, the Projected Gradient Descent (PGD) attack [8] assesses the adversarial robustness of DNNs from the robust optimization perspective.
@@
In this paper, we conduct a comprehensive empirical evaluation to assess the SDMs that DTI uses in the laptop refurbishing process regarding detection accuracy, prediction uncertainty, and adversarial robustness. Specifically, we adopt MC-Dropout as the UQ method to capture the uncertainty in model predictions. We run the model to perform multiple predictions and based on these predictions, we calculate two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We employ Dense Adversary Generation (DAG) [9] as the adversarial attack and define robustness score (RS) to measure the robustness of the SDMs. RS measures robustness from two perspectives: robustness concerning predictive precision and robustness concerning prediction uncertainty. Regarding benchmark datasets, we construct benchmark datasets from three data sources: datasets provided by our partner DTI, datasets synthesized by prompting two vision language models (VLMs), i.e., DALL-E-3 [10] and Stable Diffusion-3 [11], and datasets created using the adversarial attack, i.e., DAG. Our evaluation results show that different SDMs achieve different performance regarding different evaluation metrics. Specifically, regarding sticker detection accuracy, Faster R-CNN_v2 is recommended as the best SDM, while RetinaNet_v2 achieved the overall best performance regarding prediction uncertainty. Regarding adversarial robustness, Faster R-CNN_v2 and RetinaNet_v2 are recommended as the best SDMs. We also provide guidelines for SDM selection and lessons learned.
@@
In summary, our contributions are: 1) we conduct an empirical study to evaluate the detection accuracy, uncertainty, and robustness of six SDMs in detecting stickers on the laptop; 2) we employ MC-Dropout as the UQ method and calculate UQ metrics to measure uncertainty in classification and regression; 3) we present a novel robustness metric, i.e., RS, to measure the robustness by considering model uncertainty; 4) we provide guidelines for SDM selection from different perspectives.
@@ 2.1 Industrial Context
DTI is a leading Danish institute focused on developing advanced technology and transferring the technology to industry. A key area that DTI focuses on is robotics. Within robotics, DTI is leading an initiative for robotic applications for sustainable operations deemed critical in the context of CEAP. Adopting circular economy principles from the CEAP could substantially boost the EU’s GDP by 0.5% by 2030, creating approximately 700,000 new jobs [1]. Electrical and electronic equipment, one of the fastest-growing waste streams in the EU, currently sees less than 40% recycled, resulting in significant value loss [1].
@@
This scenario underscores the need for refurbishing electronics such as laptops with robots, which is the context of this paper. A typical refurbishment process consists of disassembly, cleaning, inspection, restoration, reassembly, and testing steps. Collaborative robots performing these steps often require human involvement for challenging tasks (e.g., non-trivial disassembly tasks). In particular, an essential and time-consuming step in the refurbishment process is cleaning the stickers on the laptop, which can be further divided into two tasks: sticker detection and removal. Accurate detection of stickers with high confidence is crucial, as incorrect detection may result in damage to the laptop surface or incomplete sticker removal. Although the sticker detection task is straightforward for humans, the unlimited number of stickers and ever-changing types of laptops pose a significant challenge for automating and scaling this process. This issue of handling extreme variability is common across many loops of the circular economy, including recycling, refurbishing, repairing, and remanufacturing. Therefore, having a method to manage this variability in a scalable and automated way is crucial for becoming more circular.
@@
Robotics and AI systems can automate the sticker detection process. To this end, DTI leveraged computer vision technology and developed DNN-based SDMs to identify stickers on laptops, turning the sticker detection task into an automated process. During the DNN designing and training process, uncertainty might arise due to inappropriately designed model architectures or insufficient training data. For example, some laptop brands may not be included in the training data, so the SDMs may not detect stickers correctly or behave highly uncertainly (e.g., detecting the same object multiple times but giving different detection results). Therefore, we must quantify and manage the uncertainty in the refurbishment process of electronic products. Specifically, quantifying uncertainty in SDMs is crucial from several perspectives. First, it ensures safe and reliable decisions of robots, especially when robots work alongside humans. Besides, managing detection uncertainty can significantly improve quality control, reduce errors, and increase accuracy. In addition, adaptability can be enhanced by helping the model adapt to different sticker types and conditions, effectively guiding human intervention when needed. Finally, workflow efficiency can be optimized by prioritizing cases that require human attention, thereby reducing downtime and streamlining workflows.
Figure 1: Overview of the Uncertainty Quantification Process for Sticker Detection Models.
@@ 2.2 Bayesian Uncertainty Quantification in Deep Learning
Gal and Ghahramani [6] proposed MC-dropout as a Bayesian approximation method, which shows that introducing dropout in DNN can be interpreted as the approximation of a probabilistic Bayesian model in deep Gaussian processes. Dropout is commonly applied in deep learning models to avoid over-fitting by randomly dropping units along with their connections from the neural network [13]. To this end, the predictive distribution is approximated by sampling multiple predictions:
$p(y^{*}\mid x^{*},\mathbf{X},\mathbf{Y})\approx\frac{1}{T}p(y^{*}\mid x^{*},\mathbf{w}^{{}^{\prime}}),$ (1)
where $x^{*}$ is the input sample, $y^{*}$ is the corresponding predicted output, $T$ is the number of predictions, and $w^{*}$ is the model weights after dropout. The $T$ DNN predictions are different because neurons are randomly dropped in each prediction, based on which we quantify the uncertainty of the overall predictions.
@@ 3 Methodology
Figure 1 presents an overview of the UQ process for the SDMs. As the figure shows, the UQ process consists of dataset creation, MC-dropout-based prediction, and UQ metrics calculation. The dataset creation constructs benchmark datasets from three data sources: datasets from DTI, datasets synthesized by prompting two VLMs, and datasets created using an adversarial attack technique. For each image in the benchmark datasets, the SDM makes T predictions based on the MC-Dropout method to capture the prediction uncertainty. We then employ the density-based algorithm, HDBSCAN, to cluster the detected stickers in T predictions and calculate the UQ metrics based on the detected stickers. We consider two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We present the MC-Dropout method in Section 3.1 and the UQ metrics in Section 3.2. Sections 3.3 and 3.4 introduce VLM-based and adversarial attack-based dataset creation.
@@ 3.1 Monte Carlo Dropout-Based Uncertainty Quantification
Bayesian probabilistic theory [14] has been shown as a universal theoretical framework for uncertainty reasoning, and as discussed in Section 2.2, MC-dropout is a Bayesian approximation for quantifying uncertainty. Hence, we adopt MC-Dropout as the UQ method and inject inference-time activated dropout layers into the pre-trained SDMs. By doing so, we build uncertainty-aware variants of the original models. Specifically, for a SDM $f$, we first build its variant as $f^{p}$ by applying dropout layers with dropout rate $p$. Next, we let $f^{p}$ make $T$ predictions for a given image input $X$ to obtain a sample of outputs $O=\{f^{p}_{t}\}_{t=1}^{T}$. The outputs $O$ represent a sample from the model’s predictive distribution, and MC-Dropout quantifies the uncertainty of $f_{p}$’s outputs by extracting information regarding the variability of the $T$ predictions. We set $T$ to 20 following the guidelines by Gal and Ghahramani [6].
@@ 3.2.3 Uncertainty Metrics for Label Classification
$\mathbf{y}=\{\operatorname*{argmax}(sm_{k})\}_{k=1}^{W},$ (2)
@@
$c^{*}=\operatorname*{argmax}_{c=1,...,C}\sum_{k}\mathds{1}[\mathbf{y}^{k}=c],$ (3)
where $\mathds{1}$ is the indicator function. Then the number of times $c^{*}$ was sampled is $\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]$. Finally, VR is calculated as:
$\textit{VR}=1-\frac{\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]}{W}.$ (4)
@@
$SE=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c)),$ (5)
@@
MI quantifies the information difference between the predicted and posterior of the model parameters, providing a different uncertainty measure for classification tasks. For the detected object in each cluster, MI can be approximated as [16]:
$\begin{split}\textit{MI}=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\\ +\frac{1}{W}\sum_{k=1}^{W}\sum_{c=1}^{nc}sm_{k}(c)\times\log sm_{k}(c),\end{split}$ (6)
@@ 3.2.4 Uncertainty Metrics for Bounding Box Regression
$\textit{TV}=\sum_{v\in box}\delta_{v}^{2}.$ (8)
TV ranges in $[0,+\infty)$, and a larger TV value indicates a higher uncertainty.
@@
$\begin{split}CP=\{(x_{1},y_{1})_{k},(x_{2},y_{2})_{k},(x_{2},y_{1})_{k},(x_{1},y_{2})_{k}\}_{k=1}^{W},\end{split}$ (9)
@@
$\textit{PS}=\frac{1}{|CP|}\sum_{cp\in CP}Area(convexh_{cp}).$ (11)
@@
Recall that for an input image $X$, multiple objects will be detected, so after performing the HDBSCAN clustering algorithm, multiple clusters will be obtained, each cluster representing a detected object. To get the uncertainty of the input image $X$, for each UQ metric, we take the average of all metric values of all detected objects in $X$. Figure 2 shows an example of using the SDM to detect stickers on a laptop and how to calculate UQ metrics based on the model outputs. The SDM makes T predictions, and each prediction returns a label based on the softmax score and a bounding box containing four corner points. The convex hull of each corner point is then calculated. The UQ metrics for label classification are then calculated based on the softmax scores in T predictions. Regarding UQ metrics for bounding box regression, TV is calculated based on the variance of the values of each corner point, and PS is calculated based on the area covered by convex hulls.

### v2

# Assessing the Uncertainty and Robustness of the Laptop Refurbishing Software

@@ Abstract
Refurbishing laptops extends their lives while contributing to reducing electronic waste, which promotes building a sustainable future. To this end, the Danish Technological Institute (DTI) focuses on the research and development of several robotic applications empowered with software, including laptop refurbishing. Cleaning represents a major step in refurbishing and involves identifying and removing stickers from laptop surfaces. Software plays a crucial role in the cleaning process. For instance, the software integrates various object detection models to identify and remove stickers from laptops automatically. However, given the diversity in types of stickers (e.g., shapes, colors, locations), identification of the stickers is highly uncertain, thereby requiring explicit quantification of uncertainty associated with the identified stickers. Such uncertainty quantification can help reduce risks in removing stickers, which, for example, could otherwise result in software faults damaging laptop surfaces. For uncertainty quantification, we adopted the Monte Carlo Dropout method to evaluate six sticker detection models (SDMs) from DTI using three datasets: the original image dataset from DTI and two datasets generated with vision language models, i.e., DALL-E-3 and Stable Diffusion-3. In addition, we presented novel robustness metrics concerning detection accuracy and uncertainty to assess the robustness of the SDMs based on adversarial datasets generated from the three datasets using a dense adversary method.
Our evaluation results show that different SDMs perform differently regarding different metrics. Based on the results, we provide SDM selection guidelines and lessons learned from various perspectives.
Keywords Uncertainty Quantification  $\cdot$ Robustness Evaluation  $\cdot$ Object Detection  $\cdot$ Deep Neural Network
@@ 1 Introduction
The European Union’s Circular Economy Action Plan (CEAP) highlights the need for sustainable operations to promote circular economy processes and encourage sustainable consumption [1]. One essential activity is refurbishing electronic devices, e.g., laptops, to extend their lives, reduce electronic waste, and provide affordable options for consumers. A critical and time-consuming step in refurbishment is removing stickers from the laptop by first identifying stickers and their locations. Manual cleaning is time-consuming and faces challenges in finding enough skilled workers, and current automation solutions are not built for this amount of variation, limiting sustainability and scalability. Thus, novel solutions to automate the processes are needed. Robotics offers a promising solution to simplify and scale up this process, increasing efficiency and reducing labor costs.
@@
The Danish Technological Institute (DTI) develops, applies, and transfers technology to industry and society. One leading area that DTI focuses on is laptop refurbishing automatically with robots, where software is a crucial part of all laptop refurbishing steps. The software responsible for the cleaning process in laptop refurbishment integrates deep neural networks (DNN)–based sticker detection models (SDMs) built by DTI for automatic sticker detection, which is the basis for successful automatic sticker removal. The SDMs are built on open-source object detection DNNs and trained using a sticker detection dataset specially designed by DTI. Due to inappropriate model architecture or insufficient training data, the design and training process may introduce uncertainties into the SDMs, making them vulnerable under certain conditions, such as adversarial attacks, noisy data, or unforeseen input distributions [2]. This vulnerability highlights the need to quantify uncertainty and evaluate the robustness of the SDMs and software they are integrated in, which is crucial for trustworthy sticker removal, as incorrect detection may damage the laptop surface or incomplete sticker removal.
@@
As a first step towards holistic uncertainty quantification (UQ) and handling in laptop refurbishing robotic software, we conduct a comprehensive empirical evaluation to assess the SDMs DTI uses regarding detection accuracy, prediction uncertainty, and adversarial robustness. Specifically, we adopt Monte Carlo Dropout (MC-Dropout) [3] as the UQ method to capture the uncertainty in model predictions. We run the model to perform multiple predictions, and based on these predictions, we calculate two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We employ Dense Adversary Generation (DAG) [4] as the adversarial attack and define robustness score (RS) to measure the robustness of the SDMs. RS measures robustness from two perspectives: robustness concerning predictive precision and robustness concerning prediction uncertainty. Regarding benchmark datasets, we construct benchmark datasets from three data sources: datasets provided by our partner DTI, datasets synthesized by prompting two vision language models (VLMs), i.e., DALL-E-3 [5] and Stable Diffusion-3 [6], and datasets created using the adversarial attack, i.e., DAG. Our evaluation results show that different SDMs achieve different performance regarding different evaluation metrics. Specifically, regarding sticker detection accuracy, Faster R-CNN_v2 is recommended as the best SDM, while RetinaNet_v2 achieved the overall best performance regarding prediction uncertainty. Regarding adversarial robustness, Faster R-CNN_v2 and RetinaNet_v2 are recommended as the best SDMs. We also provide guidelines for SDM selection and lessons learned.
@@
In summary, our contributions are: 1) we conduct an empirical study to evaluate the detection accuracy, uncertainty, and robustness of six SDMs implemented in robotic software to detect stickers on the laptop; 2) we employ MC-Dropout as the UQ method and calculate UQ metrics to measure uncertainty in classification and regression; 3) we present a novel robustness metric, i.e., RS, to measure the robustness by considering model uncertainty; 4) we provide guidelines for SDM selection from different perspectives.
@@ 2.1 Industrial Context
DTI is a leading Danish institute focused on developing advanced technology and transferring the technology to industry. A key area that DTI focuses on is robotics. Within robotics, DTI is leading an initiative for robotic applications for sustainable operations deemed critical in CEAP. Adopting circular economy principles from the CEAP could substantially boost the EU’s GDP by 0.5% by 2030, creating approximately 700,000 new jobs [1]. Electrical and electronic equipment, one of the fastest-growing waste streams in the EU, currently sees less than 40% recycled, resulting in significant value loss [1]. This scenario underscores the need to refurbish electronics such as laptops, which is the context of this paper. A typical refurbishment process consists of disassembly, cleaning, inspection, restoration, reassembly, and testing. Collaborative robots performing these steps often require human involvement for challenging tasks, e.g., non-trivial disassembly tasks. In particular, an essential and time-consuming step in the refurbishment process is cleaning the stickers on the laptop, which can be further divided into two tasks: sticker detection and removal. Accurate detection of stickers with high confidence is crucial, as incorrect detection may result in damage to the laptop surface or incomplete sticker removal. Although the sticker detection task is straightforward for humans, the unlimited number of stickers and ever-changing types of laptops pose a significant challenge for automating and scaling this process. This issue of handling extreme variability is common across many loops of the circular economy, including recycling, refurbishing, repairing, and remanufacturing. Thus, having a method to manage this variability in a scalable and automated way is crucial for becoming more circular.
@@
Robotics and AI systems can automate the sticker detection process. To this end, DTI integrates DNN-based SDMs in their robotic software to automatically identify laptop stickers. During the DNN designing and training process, uncertainty might arise due to inappropriately designed model architectures or insufficient training data. For example, some laptop brands may not be included in the training data, so the SDMs may not detect stickers correctly or behave highly uncertainly (e.g., detecting the same object multiple times but giving different detection results). Therefore, we must holistically quantify and manage the uncertainty in the entire refurbishment process implemented as software inside robots from different perspectives, such as environmental uncertainties, machine learning models employed, and humans interacting with robots. The context of this work is the software responsible for laptop sticker detection that integrates SDMs. Quantifying uncertainty in SDMs is crucial from several perspectives. First, it ensures safe and reliable decisions of robots, especially when robots work alongside humans. Besides, managing detection uncertainty can significantly improve quality control, reduce errors, and increase accuracy. In addition, adaptability can be enhanced by helping the model adapt to different sticker types and conditions, effectively guiding human intervention when needed. Finally, workflow efficiency can be optimized by prioritizing cases that require human attention, thereby reducing downtime and streamlining workflows.
@@ 2.2 Bayesian Uncertainty Quantification in Deep Learning
Gal and Ghahramani [3] proposed MC-Dropout as a Bayesian approximation method, which shows that introducing dropout in DNN can be interpreted as the approximation of a probabilistic Bayesian model in deep Gaussian processes. Dropout is commonly applied in deep learning models to avoid over-fitting by randomly dropping units along with their connections from the neural network [9]. Thus, the predictive distribution is approximated by sampling multiple predictions:
$\vspace{-6pt}p(y^{*}\mid x^{*},\mathbf{X},\mathbf{Y})\approx\frac{1}{T}p(y^{*}\mid x^{*},\mathbf{w}^{{}^{\prime}}),\vspace{3pt}$ (1)
where $x^{*}$ is the input sample, $y^{*}$ is the predicted output, $T$ is the number of predictions, and $\mathbf{w}^{{}^{\prime}}$ is the model weights after dropout. The $T$ DNN predictions are different because neurons are randomly dropped in each prediction, based on which we quantify the uncertainty of the overall predictions.
@@ 3 Methodology
Figure 1 presents an overview of the UQ process for the SDMs employed in the robotic software for laptop refurbishing. As the figure shows, the UQ process starts with creating benchmark datasets from three data sources: datasets from DTI, datasets synthesized by prompting two VLMs, and datasets created using an adversarial attack technique. For each image in the benchmark datasets, the robot’s camera captures the image, which is then processed by specialized robotic software for sticker detection. Specifically, the SDM in the robotic software makes T predictions based on the MC-Dropout method to capture the prediction uncertainty. The T predictions are then fed to the UQ component in the robotic software to calculate UQ metrics. The UQ component first employs the density-based algorithm, HDBSCAN, to cluster the detected stickers in T predictions and then calculates the UQ metrics based on the detected stickers. We consider two types of UQ metrics: label classification UQ metrics and bounding box regression UQ metrics. We present the MC-Dropout method in Section 3.1 and the UQ metrics in Section 3.2. Sections 3.3 and 3.4 introduce VLM-based and adversarial attack-based dataset creation.
Figure 1: Overview of the Uncertainty Quantification Process for Sticker Detection Models in the Robotic Software.
@@ 3.1 Monte Carlo Dropout-Based Uncertainty Quantification
Bayesian probabilistic theory [10] is a universal theoretical framework for uncertainty reasoning, and as discussed in Section 2.2, MC-dropout is a Bayesian approximation for quantifying uncertainty. Hence, we adopt MC-Dropout as the UQ method and inject inference-time activated dropout layers into the pre-trained SDMs. By doing so, we build uncertainty-aware variants of the original models. Specifically, for a SDM $f$, we first build its variant as $f^{p}$ by applying dropout layers with dropout rate $p$. Next, we let $f^{p}$ make $T$ predictions for a given image input $X$ to obtain a sample of outputs $O=\{f^{p}_{t}\}_{t=1}^{T}$. The outputs $O$ represent a sample from the model’s predictive distribution, and MC-Dropout quantifies the uncertainty of $f^{p}$’s outputs by extracting information regarding the variability of the $T$ predictions. We set $T$ to 20 according to existing guidelines [3].
@@ 3.2.3 Uncertainty Metrics for Label Classification
$\vspace{-1pt}\mathbf{y}=\{\operatorname*{argmax}(sm_{k})\}_{k=1}^{W},\vspace{-2pt}$ (2)
@@
$\vspace{-1pt}c^{*}=\operatorname*{argmax}_{c=1,...,C}\sum_{k}\mathds{1}[\mathbf{y}^{k}=c],\vspace{-1pt}$ (3)
where $\mathds{1}$ is the indicator function. The number of times $c^{*}$ was sampled is $\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]$. Finally, VR is calculated as:
$\vspace{-1pt}\textit{VR}=1-\frac{\sum_{k}\mathds{1}[\mathbf{y}^{k}=c^{*}]}{W}.$ (4)
@@
$\vspace{-1pt}SE=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c)),$ (5)
@@
MI quantifies the information difference between the predicted and posterior of the model parameters, providing a different uncertainty measure for classification tasks. For the detected object in each cluster, MI can be computed as [12]:
$\vspace{-1pt}\begin{split}\textit{MI}=-\sum_{c=1}^{nc}(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\times\log(\frac{1}{W}\sum_{k=1}^{W}sm_{k}(c))\\ +\frac{1}{W}\sum_{k=1}^{W}\sum_{c=1}^{nc}sm_{k}(c)\times\log sm_{k}(c),\end{split}$ (6)
@@ 3.2.4 Uncertainty Metrics for Bounding Box Regression
$\vspace{-1pt}\textit{TV}=\sum_{v\in box}\delta_{v}^{2}.$ (8)
TV ranges in $[0,+\infty)$, and larger TV means higher uncertainty.
@@
$\vspace{-1pt}\begin{split}CP=\{(x_{1},y_{1})_{k},(x_{2},y_{2})_{k},(x_{2},y_{1})_{k},(x_{1},y_{2})_{k}\}_{k=1}^{W},\end{split}$ (9)
@@
$\vspace{-1pt}\textit{PS}=\frac{1}{|CP|}\sum_{cp\in CP}Area(convexh_{cp}).$ (11)
@@
Recall that for an input image $X$, multiple objects will be detected, so after performing the HDBSCAN clustering algorithm, multiple clusters will be obtained, each cluster representing a detected object. To get the uncertainty of the input image $X$, for each UQ metric, we take the average of all metric values of all detected objects in $X$. Figure 2 shows an example of using the SDM to detect stickers on a laptop and how to calculate UQ metrics based on the model outputs. The SDM makes T predictions, and each prediction returns a label based on the softmax score and a bounding box containing four corner points. The convex hull of each corner point is then calculated. The UQ metrics for label classification are calculated based on the softmax scores in T predictions. Regarding UQ metrics for bounding box regression, TV is calculated based on the variance of each corner point, and PS is calculated based on the area covered by convex hulls.